OpenStudy (anonymous):

Suppose you have 80 ft. of fence to enclose a rectangular garden. the function A=40x-x squared, gives you the area of the garden in square feet where x is the width in feet . a. What width gives you the maximum gardening area? b. What is the maximum area?

OpenStudy (anonymous):

NEED HELP

OpenStudy (anonymous):

IAM FAILING

OpenStudy (anonymous):

is something wrong with your question.. I am figuring out and there is some mistake I found out in A=40x-x ???? Check it out again

OpenStudy (anonymous):

is something wrong with your question.. I am figuring out and there is some mistake I found out in A=40x-x ???? Check it out again

OpenStudy (anonymous):

A=40x-x squared is the equation

OpenStudy (anonymous):

Calm down, the maximum area will be a square so the side will be 20 set A = 80 and solve for x

OpenStudy (anonymous):

I dont knoow how to do the square thingy

OpenStudy (anonymous):

But what is A in the equation

OpenStudy (anonymous):

\[\sqrt{80}=40x-x\]

OpenStudy (anonymous):

Is it Y

OpenStudy (anonymous):

I dont get how to get x that is the problem

OpenStudy (anonymous):

Can you go step by step for me to get it

OpenStudy (anonymous):

sorry, x is width in ft, so actually the equation should be A=40(20)-20 since width should be 20

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

(40(20)-20)^2=A

OpenStudy (anonymous):

the equation is A=40x-x.. so.. it just left.. A=30x.. right... and every each side is 20 for the whole Area.... I gree with Rouault..... That is the same thing.. But.. the equation might be.. A=40x^2-x... I think...

OpenStudy (anonymous):

so would it be 400 is the maximum area as 20 is the width that gives me the maximum area

OpenStudy (anonymous):

is my anwser correct